Grade 12th passAlgebra

the altitude at which the weight of a body is only 64%of its weight on the surface of the earth is (

swayam
7 Years agoGrade 12th pass

1 Answer

Samyak Jain
7 Years ago
Let m be the mass of the body, g be the acceeleration due to gravity on earth’s surface, G is gravitational constant, M is the mass and R is the radius of the earth, then we know that
Weight of body = Gravitational force between body and earth   ...(1)
i.e. mg = GMm / R2             ...(2)
Let at height h from the earth’s surface, acceleration due to gravity becomes g1.
Then applying condition (1),  mg1 = GMm / (R + h)2        ...(3)
Divide (3) by (2) to get mg1 / mg  =  R2 / (R + h)2   or  mg1 / mg  = 1 / (1 + h/R)2  ..(4)
It is given that mg1 = 64% of mg  i.e.  mg1 / mg = 64/100  =  16/25    ..(5)
From (4) and (5),  1 / (1 + h/R)2  =  16/25   \Rightarrow   (1 + h/R)2 = 25/16
1 + h/R  =  5/4   \Rightarrow   h = R/4.